# Optimal Transmission Power Scheduling for Networked Control System under DoS Attack

### v1

@@ Abstract
Designing networked control systems that are reliable and resilient against adversarial threats, is essential for ensuring the security of cyber-physical systems. This paper addresses the communication-control co-design problem for networked control systems under denial-of-service (DoS) attacks. In the wireless channel, a transmission power scheduler periodically determines the power level for sensory data transmission. Yet DoS attacks render data packets unavailable by disrupting the communication channel. This paper co-designs the control and power scheduling laws in the presence of DoS attacks and aims to minimize the sum of regulation control performance and transmission power consumption. Both finite- and infinite-horizon discounted cost criteria are addressed, respectively. By delving into the information structure between the controller and the power scheduler under attack, the original co-design problem is divided into two subproblems that can be solved individually without compromising optimality. The optimal control is shown to be certainty equivalent, and the optimal transmission power scheduling is solved using a dynamic programming approach. Moreover, in the infinite-horizon scenario, we analyze the performance of the designed scheduling policy and develop an upper bound of the total costs. Finally, a numerical example is provided to demonstrate the theoretical results.
@@ I Introduction
Cyber-physical systems are systems that integrate sensors, controllers, and actuators to collaborate over a communication network for regulating and optimizing the behavior of a dynamic system [1, 2]. There are many applications such as robotics [3], smart grids [4], and intelligent vehicle [5]. These networked systems generally assume that data are measured and transmitted periodically to update control signals [6]. However, transmitting sensory data over a communication network is costly in general due to limited battery energy and communication channel bandwidth. This fact motivates us to co-design the control and communication strategies to ensure that the valuable sensory data is efficiently transmitted to the controller, thereby improving overall system performance.
@@
Due to increased connectivity, these cyber-physical systems are suffering from cyber threats. Two most common cyber attacks are deception attacks and denial-of-service (DoS) attacks [7]. Deception attacks degenerate the system performance by maliciously modifying the information contained in transmitted data [8, 9, 10]. DoS attacks render data packets unavailable by jamming communication channels [11, 12]. It is more destructive compared to deception attacks when facing threats from large-scale and persistent attacks. Considering the detrimental impact of cyberattacks on control performance, proactive risk management strategies are essential to ensure the resilience of networked control systems.
@@
Transmission power scheduling means that the transmission device dynamically adjusts power levels according to the changing network conditions to fulfill requirements of cyber-physical systems [13, 14]. Compared to the traditional event-based triggering, power scheduling allows for continuous optimization of energy usage, which is crucial for battery-powered devices and energy-constrained environments. Increasing transmission power levels typically enhances communication link reliability. As a consequence, the power scheduler envisions a tradeoff between energy usage and improved control performance by reliable transmission. To address this issue, the authors in [14] investigate the optimal power scheduling in remote estimation. Moreover, [13, 15] investigate the co-design of power scheduling and control strategies for networked control. However, none of the above works take cyber attacks into consideration. Some works, e.g., [16, 17], investigate the power allocation under the DoS attack to minimize the long-term mean square error covariance and propose a variance-based scheduling mechanism. However, variance-based scheduling does not utilize real-time innovations when making decisions and thus is generally outperformed by state-based scheduling [18]. In real-world applications such as financial investment, current rewards are typically valued more than future rewards. Thus, this article aims to obtain a co-design strategy that minimizes the expectation of a discounted accumulated cost.
@@
In this work, we propose a framework to jointly co-design the control law and power scheduler for the networked control system under DoS attack. We consider a signal-to-interference-plus-noise ratio (SINR)-based network model [17], where the transmission success probability is affected by the transmission power level chosen by the scheduler and the energy imposed by the attacker. The contribution of this work is summarized as follows. For both cases, the optimal co-design of control and scheduling is shown to be separable, given that the knowledge about the attack energy is symmetric between the controller and the scheduler. The optimal control law is shown to be certainty equivalent. By applying the dynamic programming approach to the remaining sequential decision problem, the condition for optimal power scheduling is derived. Further, in the infinite-horizon case, we provide an upper bound on the total cost achieved by the designed power scheduler.
@@
The remainder of this article is structured as follows: Section II introduces preliminaries. Section III and Section IV present the main result on optimal co-design of control and transmission power scheduling in finite-horizon and infinite-horizon cases, respectively. Section V presents numerical simulations. Section VI concludes this work.
@@
Notations. Denote by $\mathbb{R}$ and $\mathbb{R}^{n}$ the set of real numbers and the set of the $n$-tuples of real numbers, respectively. Denote by $\mathbb{N}$ the set of nonnegative integers. For $x,y\in\mathbb{N}$ and $x\leq y$, the set $\mathbb{N}_{[x,y]}$ denotes $\{z\in\mathbb{N}|x\leq z\leq y\}$. Denote $x_{0:k}$ as the history of state $x_{t}$ during time $t\in\mathbb{N}_{[0,k]}$. For a random variable $X$, $X\sim\mathcal{D}_{X}$ implies that $X$ is distributed according to the distribution $\mathcal{D}_{X}$. Denote $\mathbf{E}[\cdot]$, $\mathbf{E}[\cdot\mid\cdot]$ and ${\rm{cov}}[\cdot]$ as the expectation, the conditional expectation and the covariance of the random variable, respectively. If the random variable $x$ follows a normal distribution with the mean of $a$ and the covariance of $\Sigma$, we write $x\sim\mathcal{N}(a,\Sigma)$. Denote by $\mu_{\Sigma}(X)=\frac{1}{(2\pi)^{\frac{n}{2}}|\Sigma|^{\frac{1}{2}}}\exp(-\frac{(X-a)^{\top}\Sigma^{-1}(X-a)}{2})$ the probability density function of a $n$-dimension random vector $X\sim\mathcal{N}(a,\Sigma)$. Denote by $Q_{f}(x)=\frac{1}{\sqrt{2\pi}}\int_{x}^{\infty}e^{-x^{2}/2}$ the tail function of the standard Gaussian distribution. Denote the spectral radius of $A$ as $\rho(A)$.
@@ II Preliminaries
We consider the energy-constrained feedback control system over the communication network, as illustrated in Fig. 1. The scheduler periodically determines the power of transmitting sensory data, which affects packets’ transmission success probability. The controller generates the control signal based on the remote estimate.
@@ II-B Network model
    The attack energy $a_{k}\in\mathbb{R}$, for $k\in\mathbb{N}$, is a i.i.d. random process with a continuous distribution $\mathcal{D}_{a}$.
@@
    The attack energy $a_{k}\in\mathcal{S}$ with $\mathcal{S}:=[0,a_{\max}]$ and $a_{\max}$ being a positive scalar.
@@
The attack energy $a_{k}$ can be interpreted as a channel fading parameter that encompasses unpredictable variations in the wireless channel [13]. The subsequent sections will discuss control and power scheduling co-design, considering known and unknown attack energy scenarios, respectively.
Denote the transmission success as a random binary process $\delta_{k}\in\{0,1\}$, where $\delta_{k}=1$ denotes the packet is transmitted successfully, and $\delta_{k}=0$ otherwise. The packet dropout probability is affected by the transmission power and attack energy:
@@
where $\alpha$ is a communication channel parameter and $Q_{f}(\cdot)$ is the tail function of the standard Gaussian distribution. Moreover, $q_{k}\in\mathcal{Q}:=[2Q_{f}\big{(}\sqrt{\alpha p_{\max}/\sigma^{2}}\big{)},1]$. Note that a higher transmission power indicates a lower packet dropout probability and vice versa. Assume that the network induces a one-step delay. At time $k$, the packet arrives at the remote side is $z_{k}=x_{k-1}$ if $\delta_{k-1}=1$, and $z_{k}=\emptyset$ otherwise, with $z_{0}=\emptyset$. The remote estimator is given as
@@
$\min_{f,\pi}$ $\bar{\Psi}(f,\pi)=\bar{J}_{c}(f,\pi)+\lambda\bar{J}_{p}(\pi)$ (6)
with $\bar{J}_{c}(f,\pi)$ being the infinite-horizon LQG function:
$\bar{J}_{c}(f,\pi)=\mathbf{E}\big{[}\sum_{k=0}^{\infty}$ $\gamma^{k}(x_{k}^{\top}Qx_{k}+u_{k}^{\top}Ru_{k})\big{]},$ (7)
where $\gamma\in(0,1)$, and the transmission energy consumption is $\bar{J}_{p}(\pi)=\sum_{k=0}^{\infty}\gamma^{k}p_{k}$.
@@ III Finite-horizon case
In this section, we develop the solution to Problem 1. In the following, we will decompose the co-design optimization problem and design the optimal control law.
@@ III-A Optimal control
In networked control systems, the controller may induce a dual effect: 1) stabilizing the dynamic system and 2) probing the system to reduce uncertainties. This leads to a nonclassical information structure between the power scheduler and the controller. In other words, the information available to two distributed decision makers might be asymmetric. We provide the following notion to facilitate the search for the structural results of the optimization problem.
@@
The following lemma decomposes the original co-design optimization problem, assuming that the attack energy is known, and investigates which class of policies is dominating for problem (4).
@@
with the initial value $e_{0}=x_{0}-\hat{x}_{0}$. Note that $e_{k}$ comprises primitive random variables $w_{0:k-1}$ and $e_{0}$. Substituting the algebraic Riccati equation into the cost function (4), we have
@@
$+\gamma^{k}(u_{k}+L_{k}x_{k})^{\top}\Lambda_{k}(u_{k}+L_{k}x_{k})+\gamma^{k}\lambda p_{k}\big{)}\big{]}$
@@
$2(u_{k}+L_{k}\mathbf{E}[x_{k}|\mathcal{I}_{k}^{c}])^{\top}\Lambda_{k}L_{k}e_{k}+(L_{k}e_{k})^{\top}\Lambda_{k}L_{k}e_{k}]$
@@
where the second equality is from the tower property of conditional expectation, i.e., $\mathbf{E}[(u_{k}+L_{k}\mathbf{E}[x_{k}|\mathcal{I}_{k}^{c}])^{\top}\Lambda_{k}L_{k}e_{k}]=\mathbf{E}[(u_{k}+L_{k}\mathbf{E}[x_{k}|\mathcal{I}_{k}^{c}])^{\top}\Lambda_{k}L_{k}\mathbf{E}[e_{k}|\mathcal{I}_{k}^{c}]]$ and $\mathbf{E}[e_{k}|\mathcal{I}_{k}^{c}]=0$. Then the optimal controller minimizing (11) results in (8). Moreover, we have $\Psi(f,\pi)=\Psi(f,\tilde{\pi})\geq\min_{f\in\mathcal{F}}\Psi(f,\tilde{\pi})=\Psi(f^{\ast},\tilde{\pi})=\Psi(f^{\ast},\pi^{\prime})$, where $\pi^{\prime}$ depends on $\mathcal{I}_{k}^{p}$, and the first and the last equalities follow from $\tilde{\pi}_{k}(x_{0},w_{0:k-1},a_{0:k})=\pi(\mathcal{I}_{k}^{p})$. Namely, for any feasible $(f,\pi)\in\mathcal{U}$, the set $(f^{\ast},\pi^{\prime})\in\mathcal{U}$ is a dominating class of policy pairs such that $\Psi(f,\pi)\geq\Psi(f^{\ast},\pi^{\prime})$. $\blacksquare$
Lemma 1 implies that the original co-design problem can be decomposed into two subproblems, i.e., optimal power scheduling design that depends on primitive random variables and the optimal control law design, without loss of optimality.

### v2

@@ Abstract
Designing networked control systems that are reliable and resilient against adversarial threats, is essential for ensuring the security of cyber-physical systems. This paper addresses the communication-control co-design problem for networked control systems under denial-of-service (DoS) attacks. In wireless channels, a transmission power scheduler periodically determines the power level for sensory data transmission. Yet DoS attacks render data packets unavailable by disrupting the communication channel. This paper co-designs the control and power scheduling laws in the presence of DoS attacks and aims to minimize the sum of regulation control performance and transmission power consumption. Both finite- and infinite-horizon discounted cost criteria are addressed. By delving into the information structure between the controller and the power scheduler under attack, the original co-design problem is divided into two subproblems that can be solved individually without compromising optimality. The optimal control is shown to be certainty equivalent, and the optimal transmission power scheduling is solved using a dynamic programming approach. Moreover, in the infinite-horizon scenario, we analyze the performance of the designed scheduling policy and develop an upper bound of the total costs. Finally, a numerical example is provided to demonstrate the theoretical results.
@@ I Introduction
Cyber-physical systems are systems that integrate sensors, controllers, and actuators to collaborate over a communication network for regulating and optimizing the behavior of a dynamic system [1, 2]. Its applications include robotics [3], smart grids [4], and intelligent vehicle [5]. Networked systems generally assume that sensory data is measured and transmitted periodically to update control signals [6]. However, transmitting data over a communication network is generally costly due to the limited battery energy and communication channel bandwidth. This fact motivates us to co-design the control and communication strategies ensuring that the valuable sensory data is efficiently transmitted to the controller, thereby improving overall system performance.
@@
Due to increased connectivity, cyber-physical systems are suffering from cyber threats. Two most common cyber attacks are deception attacks and denial-of-service (DoS) attacks [7]. Deception attacks degenerate the system performance by maliciously modifying the information contained in transmitted data [8, 9, 10]. DoS attacks render data packets unavailable by jamming communication channels [11, 12, 13]. It is more destructive compared to deception attacks when facing threats from large-scale and persistent attacks. Considering the detrimental impact of cyberattacks on control performance, proactive risk management strategies are essential to ensure the resilience of networked control systems.
@@
Transmission power scheduling means that the transmission device dynamically adjusts power levels according to the changing network conditions to fulfill requirements of cyber-physical systems [14, 15]. Compared to the traditional event-based triggering, power scheduling allows for continuous optimization of energy usage, which is crucial for battery-powered devices and energy-constrained environments. Increasing transmission power levels typically enhances communication link reliability. As a consequence, the power scheduler envisions a tradeoff between energy usage and improved control performance. In networked control systems, power scheduling is typically addressed by selecting a power level from a predetermined finite set to achieve an optimization objective, such as [16, 17]. Another common approach involves pre-defining a specific scheduling policy and searching for the optimal scheduling parameters by solving associated optimization problems, such as [18, 19]. However, these methods often constrain the scheduling policy’s structure, leading to suboptimal solutions. In contrast, [14, 15, 20] do not fix the structure of scheduling policies. Instead, they focus on finding the optimal mapping from the system state and communication channel conditions to the power scheduling decision. More specifically, [15] investigates the optimal power scheduling strategy for remote estimation over a fading channel. Similarly, [14, 20] investigate the joint co-design of power scheduling and control, aiming to minimize the long-term transmission energy and control cost. However, the above works do not consider the impact of cyber attacks.
When addressing attacks in networked control systems, this work differs from existing works that focus on the energy allocation of the attacker [21, 22, 23] or joint energy allocation design of the attacker and the power scheduler [24]. Instead, we focus on the optimal power scheduler design under DoS attack. The relating works are [25, 26], where [25] investigate the power allocation under the DoS attack to minimize the long-term mean square error covariance and propose a variance-based scheduling mechanism. However, variance-based scheduling does not utilize real-time innovations when making decisions and thus is generally outperformed by state-based scheduling [27]. Moreover, [26] investigates a multi-channel schedule for remote estimation under DoS attack, where the power is chosen from a pre-determined level set. In contrast, our scheduling method chooses the power values from a continuous real-valued domain, which allows for greater flexibility in design. Additionally, in real-world applications, current rewards are typically valued more than future rewards. Thus, this article aims to obtain a co-design strategy that minimizes the expectation of a discounted cumulated cost.
@@
In this work, we propose a framework to jointly co-design the control law and power scheduler for the networked control system under DoS attack. We consider a signal-to-interference-plus-noise ratio (SINR)-based network model [24], where the transmission success probability is affected by the attack energy and the transmission power level chosen by the scheduler. The contribution of this work is summarized as follows. For both the finite- and infinite-horizon cases, the optimal co-design of control and scheduling is shown to be separable, given that the knowledge about the attack energy is symmetric between the controller and the scheduler. The optimal control law is shown to be certainty equivalent. We apply the dynamic programming approach to the remaining sequential decision problem. In the finite-horizon case, we provide the analytical solution of the optimal state-based power scheduling design and its greedy version that simplifies computation. In the infinite-horizon case, we solve the corresponding Bellman equation on bounded Borel state space with discounted cost and derive the optimal stationary power scheduling policy. Further, we establish an upper bound on the total cost achieved by the designed power scheduler.
@@
The remainder of this article is structured as follows: Section II introduces preliminaries. Section III and Section IV present the main result on optimal co-design of control and transmission power scheduling in finite- and infinite-horizon cases, respectively. Section V presents numerical simulations. Section VI concludes this work.
@@
Notations. Denote by $\mathbb{R}$ and $\mathbb{R}^{n}$ the set of real numbers and the set of the $n$-tuples of real numbers, respectively. Denote by $\mathbb{N}$ the set of nonnegative integers. For $x,y\in\mathbb{N}$ and $x\leq y$, the set $\mathbb{N}_{[x,y]}$ denotes $\{z\in\mathbb{N}|x\leq z\leq y\}$. Denote $x_{0:k}$ as the history of state $x_{t}$ during time $t\in\mathbb{N}_{[0,k]}$. For a random variable $X$, $X\sim\mathcal{D}$ implies that $X$ is distributed according to the distribution $\mathcal{D}$. Denote $\mathbf{E}[\cdot]$, $\mathbf{E}[\cdot\mid\cdot]$ and ${\rm{cov}}[\cdot]$ as the expectation, the conditional expectation and the covariance of the random variable, respectively. If the random variable $x$ follows a normal distribution with the mean of $a$ and the covariance of $\Sigma$, we write $x\sim\mathcal{N}(a,\Sigma)$. Denote by $\mu_{\Sigma}(X)=\frac{1}{(2\pi)^{\frac{n}{2}}|\Sigma|^{\frac{1}{2}}}\exp(-\frac{(X-a)^{\top}\Sigma^{-1}(X-a)}{2})$ the probability density function (p.d.f) of a $n$-dimension random vector $X\sim\mathcal{N}(a,\Sigma)$. Denote the spectral radius of $A$ as $\rho(A)$.
@@ II Preliminaries
We consider the energy-constrained feedback control system over the communication network, as illustrated in Fig. 1. The scheduler periodically determines the power to transmit sensory data, which affects the transmission success probability of the packets. The controller generates the control signal based on the remote estimate.
@@ II-B Network model
    The attack energy $a_{k}\in\mathbb{R}$, for $k\in\mathbb{N}$, is a i.i.d. random process with a distribution $\mathcal{D}_{a}$.
@@
    The attack energy $a_{k}\in\mathcal{S}$ with $\mathcal{S}:=[0,a_{\max}]$ and $a_{\max}$ being a nonnegative scalar.
@@
The attack energy $a_{k}$ can be interpreted as a channel fading parameter that encompasses unpredictable variations in the wireless channel [14]. It can be measured using real-time monitoring systems. For instance, telecom networks measure DoS attack energy by evaluating traffic volumes, bandwidth consumption, and processing loads on network elements. The subsequent sections will discuss control and power scheduling co-design, considering known and unknown attack energy scenarios, respectively.
Consider a random binary process $\delta_{k}\in\{0,1\}$, where $\delta_{k}=1$ denotes the packet is transmitted successfully, and $\delta_{k}=0$ otherwise. The packet dropout probability is affected by the transmission power and attack energy:
@@
where $\alpha$ is a communication channel parameter and $Q_{f}(\cdot)$ is the tail function of the standard normal distribution [24]. Moreover, $q_{k}\in\mathcal{Q}:=[2Q_{f}\big{(}\sqrt{\alpha p_{\max}/\sigma^{2}}\big{)},1]$. Note that a higher transmission power indicates a lower packet dropout probability and vice versa. Assume that the network induces a one-step delay. At time $k$, the packet arrives at the remote side is $z_{k}=x_{k-1}$ if $\delta_{k-1}=1$, and $z_{k}=\emptyset$ otherwise, with $z_{0}=\emptyset$. This assumption is widely used to facilitate the sequential decision processes between the scheduler and the controller, see [29, 20]. More detailed discussion will be provided in Lemma 1. The remote estimator is given as
@@
$\min_{f,\pi}$ $\hat{\Psi}(f,\pi)=\hat{J}_{c}(f,\pi)+\lambda\hat{J}_{p}(\pi)$ (6)
with $\hat{J}_{c}(f,\pi)$ being the infinite-horizon LQG function:
$\hat{J}_{c}(f,\pi)=\mathbf{E}\big{[}\sum_{k=0}^{\infty}$ $\gamma^{k}(x_{k}^{\top}Qx_{k}+u_{k}^{\top}Ru_{k})\big{]},$ (7)
where $\gamma\in(0,1)$, and the transmission energy consumption is $\hat{J}_{p}(\pi)=\sum_{k=0}^{\infty}\gamma^{k}p_{k}$.
@@ III Finite-horizon case
In this section, we develop the solution to Problem 1. We will decompose the co-design optimization problem and design the optimal control law in the following.
@@ III-A Optimal control
In stochastic control systems, the control action generally has a dual effect. This means that it, on the one hand, stabilizes the system; on the other hand, reduces the system uncertainty by improving the system state estimate given the knowledge of past control actions, see [30]. It is shown in [30] that the dual effect does not exist when the conditional state estimate based on available information is independent of past control actions. When addressing the state-based scheduling of networked control systems, the dual effect can be removed by letting the scheduling policy be independent of past control signals [31]. Note that the stochastic optimization problem considered involves two decision-makers: the local scheduler and the remote estimator. We first provide the following notion to facilitate the search for the structural results of the optimization problem.
@@
The following lemma identifies which class of policies is dominating for the problem (4) when the attack energy is known. Based on it, the original co-design optimization problem (4) is decomposed into two subproblems, i.e., optimal control and optimal scheduling design.
@@
with the initial value $e_{0}=x_{0}-\hat{x}_{0}$. Substituting the algebraic Riccati equation (9) into the cost function (4), we have
@@
$+\gamma^{k}(u_{k}+L_{k}x_{k})^{\top}\Lambda_{k}(u_{k}+L_{k}x_{k})+\lambda\gamma^{k}p_{k}\big{)}\big{]}$
@@
$+2(u_{k}+L_{k}\mathbf{E}[x_{k}|\mathcal{I}_{k}^{c}])^{\top}\Lambda_{k}L_{k}e_{k}+(L_{k}e_{k})^{\top}\Lambda_{k}L_{k}e_{k}]$
@@
where the second equality follows from the tower property of conditional expectation, i.e., $\mathbf{E}[(u_{k}+L_{k}\mathbf{E}[x_{k}|\mathcal{I}_{k}^{c}])^{\top}\Lambda_{k}L_{k}e_{k}]=\mathbf{E}[(u_{k}+L_{k}\mathbf{E}[x_{k}|\mathcal{I}_{k}^{c}])^{\top}\Lambda_{k}L_{k}\mathbf{E}[e_{k}|\mathcal{I}_{k}^{c}]]$ and $\mathbf{E}[e_{k}|\mathcal{I}_{k}^{c}]=0$. Similar to the proof in [33, 6], $\mathbf{E}[(L_{k}e_{k})^{\top}\Lambda_{k}L_{k}e_{k}]$ is independent of control law $f$. This also implies that the dual effect does not exist. Then the optimal controller minimizing (11) results in the certainty equivalence controller (8). Moreover, we have $\Psi(f,\pi)=\Psi(f,\tilde{\pi})\geq\min_{f\in\mathcal{F}}\Psi(f,\tilde{\pi})=\Psi(f^{\ast},\tilde{\pi})=\Psi(f^{\ast},\pi^{\prime})$, where $\pi^{\prime}$ depends on $\mathcal{I}_{k}^{p}$, and the first and the last equalities follow from $\tilde{\pi}_{k}(x_{0},w_{0:k-1},a_{0:k})=\pi(\mathcal{I}_{k}^{p})$. Namely, for any feasible $(f,\pi)\in\mathcal{U}$, the set $(f^{\ast},\pi^{\prime})\in\mathcal{U}$ is a dominating class of policy pairs such that $\Psi(f,\pi)\geq\Psi(f^{\ast},\pi^{\prime})$. $\blacksquare$
Lemma 1 implies that the original co-design problem can be decomposed into two subproblems, i.e., the optimal control law design and the optimal power scheduling design depending only on primitive random variables, without loss of optimality. Given that the attack energy is available to both the power scheduler and the remote estimator, the information pattern between the remote controller and the local scheduler is shown to be nested. Moreover, the one-step delay setting does not impose restrictions as the local side can infer the packet arrival status according to the control signal $u_{k-1}$ and control policy $f^{\ast}$. More specifically, in the one-step delay setting, decision sequences process as $\cdots\rightarrow\{u_{k-1},\delta_{k}\}\rightarrow\{u_{k},\delta_{k+1}\}\rightarrow\cdots$, which preserves the nested property. Similarly, in the delay-free case, the nested property of the information pattern can be obtained by specifying the decision sequence ordering, such as ’$\cdots\delta_{k}\rightarrow u_{k}\rightarrow\delta_{k+1}\rightarrow u_{k+1},\cdots$’, as in [33].